News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Sunday 3 October 1999 8.00am EST

During the last 24 hours Chandra passed through the 7th eclipse of the season, exited the radiation belts, conducted an ACIS CTI measurement and began a series of HRC calibration measurements of AR Lac. These measurements will continue throughout the next orbit. The HRC door was re-opened after radiation belt exit. The spacecraft systems continued to operate nominally.

The current mission schedule command load began executing at 10.30AM 10/2 and will continue until 2.30PM 10/8.





Return to Status Reports